# Advancements in Constitutive Model Calibration: Leveraging the Power of Full-Field DIC Measurements and In-Situ Load Path Selection for Reliable Parameter Inference

### v1

@@ 3.1.3 Synthetic Data Generation
To generate synthetic experimental data of the cruciform specimen, the FEA was run with the assumed-true set of model parameters, $\boldsymbol{\theta}^{true}$, reported in Table 1. The values in $\boldsymbol{\theta}^{true}$ were loosely based on a traditional calibration [50, 51] of the same material under consideration here, so that this synthetic demonstration would closely mimic the expected behavior of a physical experiment.
@@
Table 1: Material model parameter values used for synthetic data generation
The number of load steps and the displacement increment per load step were chosen to keep the maximum total strain imposed on the cruciform within the plastic regime of deformation for an aluminum alloy, away from failure [50]. The synthetic experiment was carried out through 5 load steps with each load step imposing a displacement increment of 0.25 mm per arm. Thus a load path of 5 identical load steps, i.e. pulling on the same axis for the duration of the experiment, would result in a total displacement of 1.25 mm along that axis (and a displacement of 2.50 mm when considering symmetry) and 0 mm of displacement along the alternate axis.
To aid in the description of the simulated cruciform experiment, the displacements and resulting forces from two contrasting load paths are shown in Fig. 4. The left plot shows results from a load path pulling on Axis A for 5 consecutive load steps (AAAAA), and the right plot shows an alternating load path (ABABA). The resulting forces for each directional component are plotted vs. simulation pseudotime using parameter values from Table 1, and displacements are plotted on the right axis. The 5 diamonds correspond to nodes in the load path tree, which is where measurements were collected for calibration.
Two data types were extracted from the FEA to be used as the synthetic experimental data: these were 1) the nodal displacements (which simulate full-field DIC displacement measurements) collected from 1,022 nodes in the circular region of the cruciform geometry, indicated by the green nodes in Fig. 3, and 2) the resultant forces along each axis of the cruciform. The QoIs were stored at the end of each load step, which coincide with each node of the load path tree (Fig. 2, but extended to 5 load steps). Independent and identically distributed (i.i.d.) Gaussian noise was added to the synthetic data to mimic what is seen experimentally. For the forces, a noise variance of 582.11 $\text{N}^{2}$ was used and for the nodal displacements, a noise variance of 4 µm2 was used. In the remainder of this paper, ‘observed data’ refers to data that was synthetically generated.
An example of the synthetic displacement data for the AAAAA (Figs. 1 and 2) and ABABA (Figs. 1 and 2) load paths can be found in F and G, respectively.
@@
3.2 Dimension Reduction of DIC Data
The simulated DIC measurements provide rich information content for calibration; however, the high-dimensional representation of DIC data poses a numerical challenge for calculating the EIG (Sec. 3.4.2) and likewise in the construction of the surrogate models built for the expensive finite element analysis (Sec. 3.3).
@@ 3.2 Dimension Reduction of DIC Data
PCA is a method for dimension reduction which performs a linear transformation on a data set $A\in\mathbb{R}^{u\times v}$ in order to project it onto a new lower-dimensional orthogonal basis, a.k.a. principal components, such that the maximum amount of variance in the data is preserved. This dimension reduction is accomplished through a decomposition of the original matrix $A$ into three new matrices via SVD [55] such that $A=USV^{T}$. The maximum number of principal components that can be extracted from the data matrix is determined by the rank $r$ of matrix $A$; given a full-rank matrix, $r\coloneqq\text{rank}(A)=\min\{u,v\}$.
@@
Fig. 5 shows a schematic representation of SVD for a rectangular matrix $A$, where $u<v$. The matrices $U\in\mathbb{R}^{u\times r}$ and $V^{T}\in\mathbb{R}^{r\times v}$ are orthonormal bases with columns and rows, respectively, consisting of the left and right singular vectors of $A$, and $S\in\mathbb{R}^{r\times r}$ is a diagonal matrix of singular values. The directions of the principal components are explained by the rows of $V^{T}$ and their magnitudes by the diagonal entries of $S$. The principal components are ordered according to the amount of variance they explain in the original data set from greatest to least. Thus, their relative importance is revealed, and a subset of components can be used to explain a given amount of variance in the original data set. By retaining a number $p<r$ of principal components, the data may represented in a reduced-dimensional space with very little information loss. Further details and nuances of PCA are left to [53].
@@
Figure 5: Schematic representation of singular value decomposition (SVD) of a non-square matrix $A\in\mathbb{R}^{u\times v}$ with $u<v$. SVD factorizes the matrix into three new matrices $A=USV^{T}$, with $U\in\mathbb{R}^{u\times r}$, $S\in\mathbb{R}^{r\times r}$, and $V^{T}\in\mathbb{R}^{r\times v}$, where $r$ is the rank of $A$. The columns and rows of $U$ and $V^{T}$ contain the left and right singular vectors of $A$, respectively. The singular values are contained in $S$. By retaining a number of principal components $p<r$, matrix $A$ can be expressed in a reduced-dimensional space.
@@
The schematic representation of SVD in Fig. 5 is reflective of the context in which PCA was exercised in this work. Training data for the PCA decomposition was generated by first obtaining samples of the unknown model parameters. For this task, Halton samples [56], which consist of a deterministic sequence of prime numbers to produce a space-filling design, were used. 400 training samples were generated within the bounds specified for the unknown parameters in Table 2, $\boldsymbol{\theta}^{hal}\in\mathbb{R}^{u\times D}$, where $D$ is the dimensionality of the unknown parameter vector $\boldsymbol{\theta}$ ($D=4$ in this case), and $u=400$—all other parameters were held fixed at the values recorded in Table 1. Each parameter sample was then used as input to the FE model, $g(\boldsymbol{\theta}^{hal})$, to produce corresponding output, which was the nodal displacements and resultant forces for both the $X$ and $Y$ directional components. However, PCA was performed only for the high-dimensional displacement field, not the resultant forces.
@@
The generated displacement data comprised matrix $A$, which was a data set of $u=400$ nodal displacement files which contained $v=1,022$ values each (green nodes in Fig. 3). The important first step was to make $A$ zero-mean by subtracting the mean of the data set such that, $A^{\prime}=A-A^{\text{mean}},\quad A^{\text{mean}}\in\mathbb{R}^{1\times v}$. Fig. 6 shows the PCA analysis for the $X$ directional component of displacement after a single load step applied along Axis A (which corresponds to Node 1 in Fig. 2). Note, separate PCA analyses were performed for each directional component of displacement ($X$ and $Y$) and for each node in the load path tree. A small portion of the training data set is shown in Fig. 6(a), and the retained principal components, $V^{*T}\subseteq V^{T}$, $V^{*T}\in\mathbb{R}^{p\times v}$, and singular values of $S$ from the SVD are shown in Fig. 6(b) and (c), respectively. The singular values decrease in magnitude quickly, indicating a waning level of information content in the corresponding principal components. A number of components $p=5$ was retained, thus, reducing dimensionality of the data set from $400\times 1,022$ to $400\times 5$ after performing the linear transformation $A^{\text{pca}}=A^{\prime}V^{*}\in\mathbb{R}^{u\times p}$. The retained principal components at Node 1 in the load path tree (Fig. 2) are shown for both $X$ and $Y$ in B.
@@
Figure 6: PCA uses SVD to factor a data matrix into 3 new matrices $A=USV^{T}$ by which dimension-reduction proceeds; this approach was used to reduce the full-field displacement measurements from DIC (here, synthetically generate from FEA). The first step was to generate $u$ parameter samples, $\boldsymbol{\theta}^{hal}\in\mathbb{R}^{u\times D}$, and evaluate the FE model at the sampled parameters, $g(\boldsymbol{\theta}^{hal})$. The output was collected into the data matrix, $A$, in (a). The first five principal components in $V^{T}$ were retained to produce $V^{*T}$ and are shown in (b). The corresponding singular values of the retained components in $S$ are shown with red dots in (c), and the singular values of the components that were truncated are shown with blue dots. The images are shown for the decomposition of the X directional component after a single load step along Axis A (Node 1 in Fig. 2).
@@
The original data set can be reconstructed from the retained components with $A^{\text{recon}}=A^{\text{pca}}V^{*T}+A^{\text{mean}}$. Fig. 7(a) shows an example reconstruction of one of the data sets which comprised $A$. The point-wise absolute difference between the reconstructed and original data file is shown in Fig. 7(b), which is very small relative to the magnitude of the measurements. Fig. 7(c) shows that with only 5 retained principal components, over $99.9\%$ of the information content in the original data set was preserved. Thus, retaining just the first few components successfully describes the majority of the variation in the data set and results in minimal information loss. Furthermore, there is a negligible benefit in retaining any additional components.
@@
Figure 7: The quality of the reduced-dimensional representation of the displacement field through PCA can be visualized by (a) reconstructing the displacement field with the $p=5$ retained principal components, $A^{\text{recon}}$, and (b) calculating the point-wise absolute difference between the reconstructed field and real field, $|A^{\text{recon}}-A|$. The low error seen here is expected since the 5 principal components explain (or capture) over 99.9% of the variance in the original data set (c). Thus very little information was lost with the truncation of the remaining principal components. The example reconstruction is shown for the displacement field of the X directional component after a single load step along Axis A (Node 1 in Fig. 2). Note, the displacement control was prescribed at the end of the cruciform arms, thus resulting in a slightly lower maximum displacement in the region of interest.
@@
The new orthogonal basis which is obtained can now be used to transform any new DIC file which is on the same grid as that used for training. In an experimental example, the DIC data would need to pre-processed before applying the PCA transformation in order to obtain measured values on an identical grid as that used for the PCA training. Since the same model is being used for the data generation process and model calibration in this case, no pre-processing of the data was necessary. Thus, as experimental data is collected (or simulated in this case), the PCA transformation can be used to express the high-dimensional data $A^{\text{new}}\in\mathbb{R}^{1\times v}$ in a low-dimensional space, $A^{\text{new}}V^{*}\in\mathbb{R}^{1\times p}$.
@@
Dimension-reduction of high-dimensional data with minimal information loss is a crucial component of being able to use the highly-informative DIC data for model calibration. Numerical issues in calculating the EIG (Sec. 3.4.2) due to data dimensionality were resolved by representing the data in a low-dimensional space. Furthermore, the construction of the surrogate models, discussed subsequently in Sec. 3.3, was made viable with dimension reduction.
@@ 3.3 Surrogate Model
The combined Bayesian calibration via the Laplace approximation (Sec. 3.4.1) and EIG calculation (Sec. 3.4.2) requires a minimum of $10^{4}$ model evaluations. [^2] This number may vary based on the EIG sampling strategy and other selected settings. Each run of the FEA conservatively takes 2 minutes to run, which equates to approximately 13 days to complete a single cycle of the ICC loop. Clearly, this computational demand is not feasible for the quasi-real-time setting the ICC framework is intended to operate within. Therefore, in order to cost-effectively perform calibration of the constitutive model within the ICC framework, the model evaluation must be reduced significantly, which is accomplished through the construction of a computationally inexpensive surrogate model replacement of the expensive FEA.
Fig. 8 shows the workflow for building a surrogate replacement for the FEA. The data needed to train and test the surrogate model was created by using a space-filling Halton sequence to generate parameter samples $\boldsymbol{\theta}^{hal}$ that served as input to the FEA—Fig. 8(a)—which then returned the quantities of interest (QoIs) as output—Fig. 8(b). The displacement values were then dimensionally reduced via PCA to obtain the corresponding singular values of the retained principal components—Fig. 8(c)—via the process described in Sec. 3.2. The resulting collection of input-output pairs constitute the training and test data for the surrogates—Fig. 8(d). The trained surrogate model then takes the model parameters $\boldsymbol{\theta}$ as input and returns an approximation to the displacement singular values and load; the displacement field may then be reconstructed from the singular values via $\text{disp}_{i}^{\text{recon}}=\text{disp}^{\text{PCA}}_{i}V^{*T}_{i}+\text{disp}^{\text{mean}}_{i}$.
@@
Figure 8: A workflow showing how the surrogate models were built. The first step was to (a) generate $u$ parameter samples of $D$ dimensions ($D=4$) with a space-filling Halton sequence to serve as input to the FE model, $g(\boldsymbol{\theta}^{hal})$. The FE model yielded nodal displacements and load data for the horizontal and vertical directions, $i\in[X,Y]$ (b). The displacement data was reduced via PCA in (c) to yield a reduced-dimensional representation of displacement through singular values of the $p$ retained principal components. The input parameters, loads, and nodal displacement singular values (all marked by the green dashed rectangles) created the input-output combinations used to train the surrogate models in (d). A separate surrogate model was trained for each QoI, $h\in[\text{disp}^{\text{PCA}},\text{load}]$ and directional component $i\in[X,Y]$ at each node in the load path tree in Fig. 2.
@@
Individual surrogate models were built for each QoI (displacement singular values and load values) and each directional component ($X$ and $Y$) for each node in the load path tree (besides the root node which represents the undeformed state) to approximately map constitutive model parameters to the QoIs, $\tilde{g}(\boldsymbol{\theta})\approx g(\boldsymbol{\theta})$. For the given load path tree, the total number of surrogates needed is $|\Xi|^{T}\times 2(p+1)$. In the adopted notation, $\Xi$ is the set containing the design options at each decision point (i.e. Ξ={Axis A,Axis $\Xi=\{\text{Axis A},\text{Axis B}\}$), and $|\Xi|$ is the cardinality of the set—equal to 2 in this example. $T$ is the number of load steps in the load path tree, and $p$ is the number of retained principal components. Multiplication by 2 is to account for both directional components (X and Y), and the addition of 1 accounts for the global load values. By building surrogates for each node—and inherently under the the prescribed boundary conditions described in Sec. 3.1.2 (i.e., 0.25 mm displacement increments)—, the history dependence which is inherent in elastoplastic models was accounted for, consistent with the approach taken in [27]. In reality, experimental boundary conditions may be more complex than those assumed here; this point is further discussed in Sec. 5. The surrogate model used in this work is a Gaussian process (GP) [57] with an anisotropic squared exponential kernel.
@@
With the aid of a fast surrogate approximation of the expensive high-fidelity model, parameter inference and load step selection was able to be performed in a rapid manner. The cycle time for a single load step was reduced from nearly 2 weeks to just several minutes, thus significantly alleviating the ICC computational burden.

### v3

@@ 3.1.3 Synthetic Data Generation
To generate synthetic experimental data of the cruciform specimen, the FEA was run with a set of model parameters which are referred to as the true parameters, $\boldsymbol{\theta}^{true}$, and are reported in Table 1. The values in $\boldsymbol{\theta}^{true}$ were loosely based on a traditional calibration [53, 54] of the same material under consideration here so that this synthetic demonstration would closely mimic the expected behavior of a physical experiment.
@@
Table 1: Material model parameter values used for synthetic data generation, $\boldsymbol{\theta}^{true}$.
The number of load steps and the displacement increment per load step were chosen to keep the maximum total strain imposed on the cruciform within the plastic regime of deformation for an aluminum alloy, away from failure [53]. The synthetic experiment was carried out through 5 load steps with each load step imposing a displacement increment of 0.25 mm per arm. Thus a load path of 5 identical load steps, i.e. pulling on the same axis for the duration of the experiment, would result in a total displacement of 1.25 mm along that axis (and a total displacement of 2.50 mm when considering symmetry) and 0 mm of displacement along the alternate axis.
To aid in the description of the simulated cruciform experiment, the displacements and resulting forces from two contrasting load paths are shown in Fig. 4. The left plot shows results from a load path pulling on the X axis for 5 consecutive load steps (AAAAA), and the right plot shows an alternating load path (ABABA). The resulting forces for each directional component are plotted vs. simulation pseudotime using parameter values from Table 1, and applied displacement boundary conditions are plotted on the right axis. The 5 diamonds correspond to nodes in the load path tree, which is where measurements were collected for calibration.
Two data types were extracted from the FEA to be used as the synthetic experimental data: these were 1) the nodal displacements (which simulate full-field DIC displacement measurements) collected from 1,022 nodes in the circular region of the cruciform geometry, indicated by the green nodes in Fig. 3, and 2) the resultant forces along each axis of the cruciform. The QoIs were stored at the end of each load step, which coincide with each node of the load path tree (Fig. 2, but extended to 5 load steps). Independent and identically distributed (i.i.d.) Gaussian noise was added to the synthetic data to mimic what is seen experimentally. For the forces, a noise variance of 582.11 $\text{N}^{2}$ was used, and for the nodal displacements, a noise variance of 4 µm2 was used. In the remainder of this paper, observed data refers to data that was synthetically generated. An example of the synthetic displacement data for the AAAAA and ABABA load paths can be found in Figs. S4-S7 in the Supplementary Material.
@@
3.2 Dimension Reduction of Displacement Field Data
The displacement field data provides rich information content for calibration, and when calibrating the expensive FE model, fast surrogate replacements are used to predict the model output—load and displacement—efficiently (Sec. 3.3). The high-dimensional nature of the displacement field data challenges the construction of the surrogate models, so dimension reduction is used in order to build surrogates for the expensive FEA.
@@ 3.2 Dimension Reduction of Displacement Field Data
PCA is a method for dimension reduction which performs a linear transformation on a data set $\mathcal{A}\in\mathbb{R}^{u\times v}$ in order to project it onto a new lower-dimensional orthogonal basis, a.k.a. principal components, such that the maximum amount of variance in the data is preserved. This dimension reduction is accomplished through a decomposition of the original matrix $\mathcal{A}$ into three new matrices via SVD [58] such that $\mathcal{A}=USV^{T}$. The maximum number of principal components that can be extracted from the data matrix is determined by the rank $r$ of matrix $\mathcal{A}$; given a full-rank matrix, $r\coloneqq\text{rank}(\mathcal{A})=\min\{u,v\}$.
@@
Fig. 5 shows a schematic representation of SVD for a rectangular matrix $\mathcal{A}$, where $u<v$. The matrices $U\in\mathbb{R}^{u\times r}$ and $V^{T}\in\mathbb{R}^{r\times v}$ are orthonormal bases with columns and rows, respectively, consisting of the left and right singular vectors of $\mathcal{A}$, and $S\in\mathbb{R}^{r\times r}$ is a diagonal matrix of singular values. The directions of the principal components are explained by the rows of $V^{T}$ and their magnitudes by the diagonal entries of $S$. The principal components are ordered according to the amount of variance they explain in the original data set from greatest to least. Thus, their relative importance is revealed, and a subset of components can be used to explain a given amount of variance in the original data set. By retaining a number $p<r$ of principal components, the data may be represented in a reduced-dimensional space with very little information loss. Further details and nuances of PCA are left to [56].
@@
Figure 5: Schematic representation of singular value decomposition (SVD) of a non-square matrix $\mathcal{A}\in\mathbb{R}^{u\times v}$ with $u<v$. SVD factorizes the matrix into three new matrices $\mathcal{A}=USV^{T}$, with $U\in\mathbb{R}^{u\times r}$, $S\in\mathbb{R}^{r\times r}$ and $V^{T}\in\mathbb{R}^{r\times v}$, where $r$ is the rank of $\mathcal{A}$. The columns and rows of $U$ and $V^{T}$ contain the left and right singular vectors of $\mathcal{A}$, respectively. The singular values are contained in $S$. By retaining a number of principal components $p<r$, matrix $\mathcal{A}$ can be expressed in a reduced-dimensional space.
@@
The schematic representation of SVD in Fig. 5 is reflective of the context in which PCA was exercised in this work. Training data for the PCA decomposition was generated by first obtaining samples of the unknown model parameters. For this task, Halton samples [59], which consist of a deterministic sequence of prime numbers to produce a space-filling design, were used. 400 training samples were generated within the bounds specified for the unknown parameters in Table 2, $\boldsymbol{\theta}^{hal}\in\mathbb{R}^{u\times D}$, where $D$ is the dimensionality of the unknown parameter vector $\boldsymbol{\theta}$ ($D=4$ in this case) and $u=400$—all other parameters were held fixed at the values recorded in Table 1. Each parameter sample was then used as input to the FE model, $g(\boldsymbol{\theta}^{hal})$, to produce corresponding output, which was the nodal displacements and resultant forces for both the $X$ and $Y$ directional components. However, PCA was performed only for the high-dimensional displacement field, not the resultant forces.
@@
The generated displacement data comprised matrix $\mathcal{A}$, which was a data set of $u=400$ nodal displacement files which contained $v=1,022$ values each (green nodes in Fig. 3). The important first step was to make $\mathcal{A}$ zero-mean by subtracting the mean of the data set such that, $\mathcal{A}^{\prime}=\mathcal{A}-\mathcal{A}^{\text{mean}},\quad\mathcal{A}^{\text{mean}}\in\mathbb{R}^{1\times v}$. Fig. 6 shows the PCA analysis for the $X$ directional component of displacement after a single load step applied along the X (horizontal) axis (which corresponds to Node 1 in Fig. 2). Note, separate PCA analyses were performed for each directional component of displacement ($X$ and $Y$) and for each node in the load path tree. A small portion of the training data set is shown in Fig. 6(a), and the retained principal components, $V^{*T}\subseteq V^{T}$, $V^{*T}\in\mathbb{R}^{p\times v}$, and singular values of $S$ from the SVD are shown in Fig. 6(b) and (c), respectively. The singular values decrease in magnitude quickly, indicating a waning level of information content in the corresponding principal components. A number of components $p=5$ was retained, thus, reducing dimensionality of the data set from $400\times 1,022$ to $400\times 5$ after performing the linear transformation $\mathcal{A}^{\text{pca}}=\mathcal{A}^{\prime}V^{*}\in\mathbb{R}^{u\times p}$. The interested reader may refer to Figs. S2 and S3 in the Supplementary Material to view the retained principal components at Node 1 in the load path tree.
@@
Figure 6: PCA uses SVD to factor a data matrix into three new matrices $\mathcal{A}=USV^{T}$ by which dimension-reduction proceeds; this approach was used to reduce the full-field displacement measurements from the FE model. The first step was to generate $u$ parameter samples, $\boldsymbol{\theta}^{hal}\in\mathbb{R}^{u\times D}$, and evaluate the FE model at the sampled parameters, $g(\boldsymbol{\theta}^{hal})$. The output was collected into the data matrix, $\mathcal{A}$, in (a). The first five principal components in $V^{T}$ were retained to produce $V^{*T}$ and are shown in (b). The corresponding singular values of the retained components in $S$ are shown with red dots in (c), and the singular values of the components that were truncated are shown with blue dots. The images are shown for the decomposition of the X directional component after a single load step along Axis A (Node 1 in Fig. 2).
@@
The original data set can be reconstructed from the retained components with $\mathcal{A}^{\text{recon}}=\mathcal{A}^{\text{pca}}V^{*T}+\mathcal{A}^{\text{mean}}\approx\mathcal{A}$. Fig. 7(a) shows an example reconstruction of one of the data sets which comprised $\mathcal{A}$. The point-wise difference between the original and reconstructed data is shown in Fig. 7(b), which is very small relative to the magnitude of the measurements. Fig. 7(c) shows that with only 5 retained principal components, over $99.9\%$ of the information content in the original data set was preserved. Thus, retaining just the first few components successfully describes the majority of the variation in the data set and results in minimal information loss. Furthermore, there is a negligible benefit in retaining any additional components.
@@
Figure 7: The quality of the reduced-dimensional representation of the displacement field through PCA can be visualized by (a) reconstructing the displacement field with the $p=5$ retained principal components, $\mathcal{A}^{\text{recon}}$ and (b) calculating the point-wise difference between the real field and reconstructed field, $\mathcal{A}-\mathcal{A}^{\text{recon}}$. The low error seen here is expected since the 5 principal components explain (or capture) over 99.9% of the variance in the original data set (c). Thus, very little information was lost with the truncation of the remaining principal components. The example reconstruction is shown for the displacement field of the X directional component after a single load step along Axis A (Node 1 in Fig. 2). Note, the displacement control was prescribed at the end of the cruciform arms, thus resulting in a slightly lower maximum displacement in the region of interest.
@@
This work focuses on calibration in the spatial domain, which requires reconstruction of the displacement fields from the surrogate predictions of the PCA singular values as previously described. However, calibration may also be performed in the spectral domain, in which case the surrogate predictions of the PCA singular values would be used directly in the calibration along with dimensionally reduced experimental data. The steps needed to transform the measured DIC data are briefly discussed here. The new orthogonal basis which is obtained can be used to transform any new DIC data which is on the same grid as that used for training. In an experimental example, the DIC data would need to pre-processed before applying the PCA transformation in order to obtain measured values on an identical grid as that used for the PCA training. If the same model were being used for the data generation process and model calibration—as in this case, no pre-processing of the data would be necessary. Thus, as experimental data is collected (or simulated), the PCA transformation can be used to express the high-dimensional data $\mathcal{A}^{\text{new}}\in\mathbb{R}^{1\times v}$ in a low-dimensional space, $\mathcal{A}^{\text{new}}V^{*}\in\mathbb{R}^{1\times p}$. Further considerations of the two approaches for calibration—field vs spectral domain—can be found in Sec. 5.
@@
Dimension-reduction of the high-dimensional displacement fields with minimal information loss is a crucial component of being able to use the highly-informative DIC data for model calibration. With the use of PCA, the construction of the surrogate models for the displacement field, discussed subsequently in Sec. 3.3, was made viable, enabling calibration with rich displacement fields.
@@ 3.3 Surrogate Model
Surrogate models have proven to be invaluable in model calibration within applied mechanics and engineering, particularly for mitigating the computational cost of high-fidelity simulations [60, 61]. By providing efficient approximations of complex constitutive models, surrogate-based approaches facilitate efficient model calibration and Expected Information Gain (EIG) calculations in this work. Their ability to accelerate computations without sacrificing accuracy ensures that these tasks can be performed reliably within a significantly reduced time-frame, making them an essential tool in the ICC framework.
@@
The combined Bayesian calibration via the Laplace approximation (Sec. 3.4.1) and EIG calculation (Sec. 3.4.2) requires a minimum of $10^{4}$ model evaluations. [^1] This number may vary based on the EIG sampling strategy and other selected settings. Each run of the FEA conservatively takes 30 minutes; clearly, this computational demand is not feasible for the quasi-real-time setting the ICC framework is intended to operate within. Therefore, in order to cost-effectively perform calibration of the constitutive model within the ICC framework, the model evaluation time must be reduced significantly, which is accomplished through the construction of a computationally inexpensive surrogate model replacement of the expensive FEA.
Surrogate models were built to approximately map constitutive model parameters to the FE model output $\tilde{g}(\boldsymbol{\theta})\approx g(\boldsymbol{\theta})$ at each node in the load path tree (besides the root node which represents the undeformed state). Fig. 8 shows the workflow for building a surrogate replacement for the FEA at a single node. The solid arrows connect the steps taken to prepare the data for surrogate training, and the dashed arrows connect the surrogate training data, enclosed with green boxes, to the training step. A Halton sequence was used to generate 400 sets of model inputs $\boldsymbol{\theta}^{hal}$–Fig. 8(a)–which then produced 400 corresponding outputs from the FEA for the given node—Fig. 8(b). The outputs are the full-field displacements and load values for the $X$ and $Y$ directional components, $g(\boldsymbol{\theta}^{hal})=[\text{disp}_{i},\text{load}_{i}],\,i\in[X,Y]$. The displacement values were then dimensionally reduced via PCA to obtain the corresponding singular values of the retained principal components—Fig. 8(c)—via the process described in Sec. 3.2. The resulting collection of input-output pairs constitute the training data for the surrogates, $\mathcal{D}_{hi}=(\boldsymbol{\theta}^{hal},\mathcal{G}_{hi})$, where $\mathcal{G}=[\text{disp}_{i}^{\text{PCA}},\text{load}_{i}],\,i\in[X,Y]$ and $h\in[\text{disp}^{\text{PCA}},\text{load}]$—Fig. 8(d). The subscript $hi$ is to indicate that a separate surrogate model was constructed for each QoI and directional component.
Each trained surrogate model then takes the model parameters $\boldsymbol{\theta}$ as input and returns either an approximation to the displacement singular values or load for one of the directional components. Thus, in order to approximate the full FE model output at a given node in the load path tree, a collection of surrogates is needed, $\tilde{g}(\boldsymbol{\theta})_{hi},h\in[\text{disp}^{\text{PCA}},\text{load}],i\in[X,Y]$. The displacement field is then reconstructed from the singular values via $\text{disp}_{i}^{\text{recon}}=\text{disp}^{\text{PCA}}_{i}V^{*T}_{i}+\text{disp}^{\text{mean}}_{i}$ before performing inference and calculating the EIG. Note, whereas the experimental design parameters (i.e. load path in this case) along with the model parameters are traditionally input to surrogate models in OED applications, here, the design parameters are omitted since a separate surrogate model is built for each node in the load path tree. Therefore, obtaining a surrogate output for a specific design, or load path, corresponds to selecting a combination of surrogate models corresponding to the nodes in that load path. For example, a surrogate approximation for Path 1 in Fig. 2 is obtained by collecting the output from the surrogate models built for Nodes 1 and 3.
@@
Figure 8: A workflow showing how the surrogate models were built. The solid arrows show the steps taken to prepare the surrogate training data, and the dashed arrows connect the surrogate training data (enclosed with green boxes) to the training step. The first step was to (a) generate $u$ parameter samples of $D$ dimensions ($D=4$) with a space-filling Halton sequence to serve as input to the FE model, $g(\boldsymbol{\theta}^{hal})$. The FE model yielded nodal displacements and load data for the horizontal and vertical directional components, $i\in[X,Y]$ (b). The displacement data was reduced via PCA in (c) to yield a reduced-dimensional representation of displacement through singular values of the $p$ retained principal components. The input parameters, loads and nodal displacement singular values created the input-output pairs in $\mathcal{D}$ used to train the surrogate models in (d). A separate surrogate model was trained for each QoI, $h\in[\text{disp}^{\text{PCA}},\text{load}]$ and directional component $i\in[X,Y]$ at each node in the load path tree in Fig. 2. The surrogate predictions of the PCA singular values were used to reconstruct the displacement fields before performing inference and calculating the EIG according to the procedure described in Sec. 3.2.
@@
For the given load path tree, the total number of surrogates needed is $|\Xi|^{\mathcal{T}}\times 2(p+1)$. In the adopted notation, $\Xi$ is the set containing the design options at each decision point (i.e. $\Xi=\{\text{A},\text{B}\}$), and $|\Xi|$ is the cardinality of the set—equal to 2 in this example. $\mathcal{T}$ is the number of load steps in the load path tree ($\mathcal{T}=2$ in the abbreviated load path tree illustrated in Fig. 2 and $\mathcal{T}=5$ for the full load path tree employed here), and $p$ is the number of retained principal components ($p=5$ in this example). Multiplication by 2 is to account for both directional components (X and Y), and the addition of 1 accounts for the global load values. By building surrogates for each node under the prescribed boundary conditions described in Sec. 3.1.2 (i.e. 0.25 mm displacement increments), the history dependence which is inherent in elastoplastic models was accounted for, consistent with the approach taken in [30]. In reality, experimental boundary conditions may be more complex than those assumed here; this point is further discussed in Sec. 5. The surrogate model used in this work is a Gaussian process (GP) [62] with an anisotropic squared exponential kernel.
@@
The surrogates generally exhibited low error across all QoIs and load steps when evaluated against data reserved for testing. The absolute error for the predicted load was less than 0.003% of the true load for all load steps and directional components, and the absolute error for the predicted displacement field (which includes both the PCA + surrogate error) was less than 0.1% of the test fields. Parity plots of the predicted load, PCA singular values and displacements can be found in A in Figs. 1 & 2. Thus, with the aid of a fast surrogate approximation of the expensive high-fidelity model, parameter inference and load step selection were able to be performed in a rapid manner. The cycle time for a single load step was reduced from months to just several minutes, thus significantly alleviating the ICC computational burden.